40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Mar 29 15:47:58 2019, gautam, Update, AUX, AUX laser fiber moved from AS table to PSL table 
    Reply  Sun Mar 31 18:39:45 2019, Anjali, Update, AUX, AUX laser fiber moved from AS table to PSL table Schematic_of_experimental_setup_for_frequency_stabilisation_of_1_micron_source.png1_micron_setup_for_frequency_noise_measurement.JPG1_micron_setup_for_frequency_noise_measurement_power_levels.png
Message ID: 14501     Entry time: Fri Mar 29 15:47:58 2019     Reply to this: 14504
Author: gautam 
Type: Update 
Category: AUX 
Subject: AUX laser fiber moved from AS table to PSL table 

[anjali, gautam]

To facilitate the 1um MZ frequency stabilization project, I decided that the AUX laser was a better candidate than any of the other 3 active NPROs in the lab as (i) it is already coupled into a ~60m long fiber, (ii) the PSL table has the most room available to set up the readout optics for the delayed/non-delayed beams and (iii) this way I can keep working on the IR ALS system in parallel. So we moved the end of the fiber from the AS table to the SE corner of the PSL table. None of the optics mode-matching the AUX beam to the interferometer were touched, and we do not anticipate disturbing the input coupling into the fiber either, so it should be possible to recover the AUX beam injection into the IFO relatively easily.

Anjali is going to post detailed photos, beam layout, and her proposed layout/MM solutions later today. The plan is to use free space components for everything except the fiber delay line, as we have these available readily. It is not necessarily the most low-noise option, but for a first pass, maybe this is sufficient and we can start building up a noise budget and identify possible improvements.

The AUX laser remians in STANDBY mode for now. HEPA was turned up while working at the PSL table, and remains on high while Anjali works on the layout.

ELOG V3.1.3-